OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=6, b=2, c=?

hartnn (hartnn):

According to pythagoras, the hypotenuse is given by \(\huge c^2=a^2+b^2\)

OpenStudy (anonymous):

i got 36+4=40

hartnn (hartnn):

you got \(\huge c^2=40, \\ \huge c= \sqrt{40}\) got this ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

and that can be simplified to .. ?

OpenStudy (anonymous):

2 square root 10 ??

hartnn (hartnn):

absolutely correct! good :)

OpenStudy (anonymous):

Thank you :)

hartnn (hartnn):

welcome ^_^